OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find g(20) + f(6).

OpenStudy (texaschic101):

f(x) = x^2 + 1 f(6) = 6^2 + 1 f(6) = 36 + 1 f(6) = 37 g(x) = 3x + 1 g(20) = 3(20) + 1 g(20) = 60 + 1 g(20) = 61 g(20) + f(6) = 61 + 37 = 98

OpenStudy (anonymous):

thank you this really helped I use this to check my answers and you came up with the same one and the same way I did it so I think that is right

OpenStudy (texaschic101):

Glad to help :)